OpenStudy (hockeygirl10):

What percent of 225 is 99?

hero (hero):

I'll help you get started: The general formula for these is: \[\dfrac{part}{whole} = \dfrac{\%}{100}\] After inputting the given values you have \[\dfrac{99}{225} = \dfrac{\%}{100}\] The percent is what we need to find so label that x: \[\dfrac{99}{225} = \dfrac{x}{100}\] Now cross multiply and solve for x. Let me know if you have any questions.